How long for $1000 to grow to $1700 at 9% (both quarter and continuously)?
Answer by reviewermath(1029)

Q:
How long for $1000 to grow to $1700 at 9% (both quarter and continuously)?
A:
Solution if compounded quarterly:
1000(1 + .09/4)^(4t) = 1700
Divide both sides by 1000
(1 + .09/4)^(4t) = 1.7
Get the natural logarithm of both sides
4t*ln(1.0225) = ln(1.7)
t = ln(1.7)/(4*ln(1.0225)) = 5.96 years if compounded quarterly
Solution if compounded continuously:
1000*e^(0.09t) = 1700
Divide both sides by 1000
e^(0.09t) = 1.7
Get the natural logarithm of both sides
0.09t = ln(1.7)
t = (ln(1.7))/0.09 = 5.9 years if compounded continuously